OpenStudy (anonymous):

evaluate the integral (tan3x)dx i got: -1/3log(cos3x)

OpenStudy (apoorvk):

You got it right!! (don't forget to include the constant of integration though)

OpenStudy (anonymous):

integral tan3x dx d(3x)=3dx integral tan3x dx =integral(1/3) tan3x d(3x) tan = sin/cos integral tanx dx = integral(1/cosx) ( -d(cosx) ) =-lncosx+C integral tan3x dx =integral(1/3) tan3x d(3x) =(1/3)(-lncos3x)+C you are right :D

OpenStudy (anonymous):

it should be ln instead of log and you shoud have the absolute symbow to make sure than its not negatine sfter the ln and you should plus C at the end :D